Henfold Winter Team Blog

As the last of 2014’s regular Wednesday evening Summer Meetings and Hive Inspections approached, planning by Henfold Apiary Manager, Neil Hallam, commenced for a team of volunteers to tackle the supervision of all the colonies at Henfold over the Winter months.

The aim being to ensure that the maximum number of colonies possible would be in a good, strong and healthy condition by the commencement of the 2015 Summer season.

Neil’s blog records preparations for and the many activities undertaken by his team since September last year. Those activities will continue until our first Summer meeting on 1st April this year.

One key activity has been the regular hefting of all hives during the winter months. As the data was gathered, it was (and continues to be) recorded by Neil on a chart, below.

Hive Weight (kg) Records

Hive Weight (kg) Records (click chart to enlarge)

Thanks go to the Winter Team: Sally Bennett, Jim Cooper, Simon Ford, Robert Harding, Mike Hill, Duncan King, Mike Lowe, Pete Scroggs, Fiona Scott and Richard Woodhouse, and of course their team leader … Neil Hallam.

Neil’s Blog

14th Mar 2015

The final push on Saturday made sure we are now ready for the new season.
  • Varroa levels in each colony were estimated by counting natural fall and showed very low levels in all colonies. Our autumn and spring treatments seem to have paid off.
  • The clipboard folders for each stand now have pages for noting varroa levels, each containing Saturday’s count along with the count from last autumn.
  • Each file now also contains the appropriate extract from our winter report showing the history of the colonies and how they fared over winter and, most importantly perhaps, the next steps to ready them for teaching this season. These steps, for example taking nucleus colonies and transferring them to full size hives, cannot be done safely until temperatures allow.
  • The files for stands designated for Basic Examination preparation now each have a copy of the syllabus.
  • Henfold March 14 2015

    Scrubbed up well – click on image to enlarge

    The biggest (and messiest) task on Saturday, was cleaning and sterilising the rest of the emptied frames from last year’s activities. We had around a hundred frames to process. But with the help of water boilers kindly lent by members of the winter team in which we prepared hot washing soda solution, we made it …

  • Next year our aim has to be to do this as we progress and keep up with cleaning, sterilising and re-waxing frames as they are removed.
The next stage this year will be to rebuild the frames with foundation ready for reuse and will start at the Henfold Spring Tidy-Up day next Saturday, 21st March.
During the Tidy-Up day, Sally Bennett and I will be giving guided tours of the Apiary and presenting our safety and beekeeping procedures to new beekeepers from the Winter Course. Simon Ford will be leading the frame making and will need as many hands to help as he can get; as well as giving the new beekeepers a chance to have a go.  Can you help on the day? (Click Here)
So, Henfold is ready when you are for a great new season of beekeeping.
Neil and the Winter Team

7th Mar 2015

When we left, the bees were active, spring is springing....

When we left, the bees were active, spring is springing….

Over the past two weeks the winter team has fed the bees, taken off woodpecker and mouse guards and is now preparing for the new season start.

The next work session will be on Saturday March 14th. The main activity then will be to ‘boil up’ as many of the dirty frames we have to get them ready for re-waxing. Other final tidy up jobs will fill the day.

2nd Feb 2015 

Planning for the 14th Feb session: This will mainly consist of hefting and feeding if necessary, reconfiguring the hives ready for next season, checking equipment inventories for each stand and some maintenance work. Please let me know if you can join me then, lots to do.

24th January (Saturday)

Met with Richard Stuart and Andrew Cornwall today at Henfold to discuss what we need to prepare for next season. Andrew confirmed the requirements for 12 training hives in the ‘circle’ and 3 intermediate set ups in the bottom copse. This means we can now start positioning hives and planning for the spring.

Whilst at Henfold, the hives were hefted. Results generally were of a slight upturn in hive weights. This could be measurement accuracy, different hefter, or environmental changes, but does not give cause for concern or indicate need for feeding yet.

I will shortly set up the next team visit to further our winter work and preparations for next season.

17th January (Saturday)

 We achieved a lot today:

  • Hefted the hive to assess stores consumption.
  • This showed the reduction in stores with the weather turning cold. Our conclusion was that whilst there was no immediate need for feeding, if the cold weather continues feeding may become advisable in February.Furtherhefts will tell. Feeding fondant sugar and pollen (Neopoll) is planned for the
    Oxalic Acid solution applied with a 'trickle dispenser' to kill varroa mites in the absence of brood.

    Oxalic Acid solution applied with a ‘trickle dispenser’ to kill varroa mites in the absence of brood. (click image to enlarge)

    end of February to kick-start the season.

  • Checked varroa levels in each colony by natural fall. The highest infestation with varroa proved to be hive J with 7300 varroa (Beebase calculated estimate of total in hive); 4 hives had less than 500; the rest over 1000. We decided to treat all colonies with Oxalic acid solution. Veterinary records were created following the administration of the treatments. A further varroa count will be made in February to check the efficacy of this treatment.
  • Where practicable, woodpecker attack prevention was applied to each of the hives. This consisted of covering exposed wood surfaces with thick plastic sheeting to try to prevent woodpeckers gaining footholds in order to drill the wood for honey and bees.
  • Safety stays have been added to the solar wax melters, completing modifications to the lids, which have included installing hinges, handles and sealing the glass, to try to prevent the lids being left ajar as was a problem a couple of times last year. Next we will be spring cleaning the metalwork ready for next season.
  • We have marked out the positions of stands for next year and in February will reposition hives to this pattern. This will provide 12 teaching positions in the ‘circle’ and 3 in the copse, one of which will have room for a circle of nucs for use in queen rearing.
  • Thanks again team for a great effort this morning, much appreciated.

Next time (I’ll e-mail dates etc as usual), amongst others, I would like us to start repositioning hives while the bees are clustered, clean up the metalwork in the wax melters, check stand heights and build up if necessary, heft and prepare for spring feeding, and (time permitting) start an inventory verification process.

7th January (Wednesday)

 Varroa boards in. Hefting planned on  Saturday 17th, 10:30 start.

16th December (Tuesday)

The figures from the latest heft of the hives shows small changes again with the exception of a couple we will need to keep an eye on,

Work started on the solar wax melters, the aim being to alter the lids so they can’t be left ajar so easily.

The next Winter Team visit to the apiary will now be after the holidays – I’ll put out an invitation.

The agenda will be to complete the wax melter modifications, heft and check the hives and plan the movement of hives ready for the next season of teaching. I’ve spoken to Andrew Buchanan and Roy Cottington separately and a provisional idea of what we might do is developing, but Andrew is going to get together with Roy and I to finalise groundwork and moving plans.

In the meantime, all the best to you and yours for the Holidays.

21st November (Friday)

Thank you for your help this morning checking the hives. We hefted the hives and generally saw only slight changes. The three sets of data gathered so far look like this:

The dramatic weight losses in two hives between the first and second hefting was due to feeder removal and further feeding may have resulted in some of the initial increases. Between the second and current hefting, hardware was changed on hive J, otherwise reduction in bee numbers and change in stores are the main factors.

Next visit we plan to modify the solar wax melter lids so they seal better, heft again and continue other maintenance.

The Henfold Record book (containing hive and Veterinary Medicine records) has been left in the container with the hefting scale.

14th November

Saturday’s session (15th November) is now cancelled due to mud. Apparently Henfold is impassable because of the building works so we have to postpone.

18th October (Saturday)

Another great team effort today.

Activities include:

  • Checked hives, removing remaining feeders.
  • Cleaned feeders and put into storage.
  • Carried out second ‘electronic’ hefting, mainly showing small decreases in weight, possibly due to more activity in the mild weather meaning a net use of stores.
  • Continued  scraping and hot soda solution clean up of frames. To date this has recovered 77 brood frames and 45 super frames for re-use.
  • Continued general container tidy up.

5th October (Sunday)

Fantastic team effort on Saturday.

Activities included:

  • Checked hives, removing feeders where possible (four colonies still a little light, two now with rapid feeders, two with fondant)
  • Using luggage scales to heft (weigh) the hives.

    Using luggage scales to heft (weigh) the hives.
    (click image to enlarge)

    Cleaned feeders and put into storage.

  • Modified hive stands for hefting
  • Carried out first ‘electronic’ hefting giving baseline for monitoring.
  • Made excellent start to scraping and hot soda solution clean up of frames (established method and a great start to recovering the 100 or so dirty frames left in the container for re-use).
  • Carried out general container tidy up.

3rd September (Wednesday)

Activities included:

  • Completed installation of mouse guards
  • Fed two Nucs
  • Applied second Apiguard doses to 4 colonies
  • Cleaned out the solar wax melters
  • Started Veterinary Medicine Administration Records for the Apiary
  • Started process of drying out the smoker fuel
  • Found handover records for hives G, J and Jnuc.

31st August (Sunday)

Activities included:

  • We fed the colonies at Henfold this afternoon and replaced 6 contact feeders with new rapid feeders.
  • Hives G and J are still to be ‘handed over’ – I will arrange this on Wednesday.
  • On Wednesday (3rd September) we need to add second Apiguard doses to 4 colonies. I will be at Henfold at around 5 to do this – let me know if you can make it.

On September 10th, 5 other colonies need Apiguard doses. Let me know your availability on the 10th or on Saturday 13th to do this and some other tidy up jobs.

August – Winter Team Set Up

To all volunteers – Can we get together briefly after announcements on Wednesday, please? Just to meet each other and so I can introduce the process of the mentors handing over their hives to us. The form I’m asking them to use and will have copies on each hive for them to start completing on Wednesday.

Leave a Reply